OpenStudy (anonymous):

Will give medal and fan. Solve for x: ln(x+9)-ln(x)=1

OpenStudy (haichi):

I dont understand this 1 soz

OpenStudy (anonymous):

Oops, ln(x+9)-ln(x)=1

OpenStudy (dumbcow):

property of logs \[\log a - \log b = \log (\frac{a}{b})\]

OpenStudy (anonymous):

ln(x+9/x)=1

OpenStudy (anonymous):

ln(x+9)-ln(x)=1 ln((x+9)/x)=1 (Log Laws) (x+9)/x=e x+9=ex 9=(e-1)x x=9/(e-1)

OpenStudy (anonymous):

How did you get from x+9=ex to 9=(e-1)x

OpenStudy (anonymous):

x+9=ex bring the x to the other side 9=ex-x factor out the x 9=(e-1)*x